An airplane flew for 4 hours with a 15mph tail wind. The return flight against the same wind took 5 hours. Find the speed of the airplane still in air.
Answer by mananth(16949)

let plane speed = x mph
d=rt
distance is same both ways
4(x+15)=5(x-15)
4x+60=5x-75
5x-4x=75+60
x=135 mph